# a possible strange quark star using the Cherenkov Telescope Array Observatory

### v1

@@ Abstract
The hypothesis that supernova remnants are key sources of Galactic cosmic rays gains support from evidence that HESS J1731-347 one of the few Galactic objects capable of accelerating hadronic cosmic rays to TeV energies may harbor an exotic strange quark star rather than a conventional neutron star. This conclusion stems from its unusually low mass and compact radius, which challenge standard neutron star models. If confirmed, such a quark star could generate cosmic rays through the transition from the two-flavor color-superconducting (2SC) phase to the color-flavor-locked (CFL) phase, potentially releasing strangelets, hypothetical strange quark matter particles. Detecting these strangelets in cosmic rays would provide groundbreaking evidence for quark matter. The future Cherenkov Telescope Array (CTA), with its unmatched sensitivity and spectral resolution in the very-high-energy (VHE) gamma-ray regime, is uniquely positioned to search for their annihilation or decay signatures. We analyze theoretical predictions for these gamma-ray signals and evaluate CTA’s potential to detect or constrain them. Additionally, we present an in-depth assessment of CTA observations of HESS J1731-347, focusing on spectral features that could confirm strangelet production. A positive detection would not only validate the existence of strange quark stars but also establish a direct link between quark matter and cosmic-ray acceleration, reshaping our understanding of compact objects and high-energy astrophysics.
@@ 1 Introduction
Should strangelets undergo disintegration or annihilation, they could produce distinctive spectral features in the gamma-ray spectrum. The Cherenkov Telescope Array Observatory (CTA), with its exceptional sensitivity and energy resolution, is ideally suited to search for these signals. As the most advanced facility for gamma-ray astronomy, the CTA is designed to detect very-high-energy (VHE) gamma rays in the energy range of 20 GeV to over 300 TeV. It can probe extreme astrophysical environments, including active galactic nuclei, supernova remnants, and potentially exotic phenomena involving strangelets or dark matter. With superior sensitivity, resolution, and sky coverage compared to existing instruments like MAGIC, VERITAS, and H.E.S.S., the CTA is poised to investigate high-energy astrophysical phenomena, such as possible signatures of strangelets in the context of HESS J1731-347. This document summarizes the theoretical predictions for strangelet-related gamma-ray signatures, their detectability with the CTA, and their implications for constraining the properties of strangelets and HESS J1731-347.
@@ 2.1 Strange Quark Matter and Current Searches
The concept of Strange Quark Matter (SQM) was first proposed in the 1980s as a distinct state of hadronic matter, coexisting alongside conventional nuclear matter Bodmer (1971); Witten (1984); Farhi and Jaffe (1984). The mass of strangelets within compact stars can range from a minimum stable mass of approximately $10sA$ Chodos et al. (1974) to as high as $A\simeq 10^{57}$, equivalent to a solar mass of approximately $2\times 10^{30}$ kg Haensel et al. (1986); Madsen (2005). SQM could have formed in the early universe shortly after the Big Bang Chodos et al. (1974), potentially contributing to baryonic dark matter Atreya et al. (2014). It may reside within the cores of neutron stars or exist as ‘‘strange quark stars,’’ which could be either purely composed of quarks Haensel et al. (1986); Alcock et al. (1986) or a mixture of hadrons and quarks Drago et al. (2016); Drago and Pagliara (2016); Drago et al. (2014). Strangelets could also be accelerated in the ergosphere of black holes Bianchi et al. (2020) or produced in collisions between binary neutron stars Bucciantini et al. (2022); Drago and Pagliara (2018); Wiktorowicz et al. (2017), potentially reaching Earth De Rújula and Glashow (1984).
Despite extensive searches, no definitive evidence for strangelets has been found, though experimental efforts have established upper limits on their production. The CERN NA52 experiment, which investigated strangelet production in Pb-Pb collisions at 158 GeV/c per nucleon, found no evidence and set upper limits on production yields for strangelets with mass-to-charge ratios up to $m/|Z|\approx 60$ GeV/c2 Appelquist et al. (1996). Similarly, the PAMELA experiment, which searched for cosmic-ray strangelets from 2006 to 2009, detected no candidates and established an upper limit on the strangelet flux for particles with charges $1\leq Z\leq 8$ and baryon numbers $4\leq A\leq 1.2\times 10^{5}$ Lewis (2005).
@@ 2.2 HESS J1731-347 and Strong First-Order (Deconfinement) Phase Transition
The High Energy Stereoscopic System (H.E.S.S.), an array of imaging atmospheric Cherenkov telescopes in Namibia, discovered the very-high-energy (VHE) gamma-ray source HESS J1731-347. Located in the constellation Scorpius near the Galactic plane, it has celestial coordinates of right ascension (RA) 17h 31m, declination (Dec) $-34^{\circ}47^{\prime}$, and galactic coordinates of longitude $l\simeq 353.6^{\circ}$ (near the galactic center) and latitude $b\simeq-0.7^{\circ}$ (slightly below the galactic plane). Its spatial coincidence with the shell-type supernova remnant (SNR) G353.6-0.7 suggests a potential association. The gamma-ray emission, observed at energies above 100 GeV, may result from accelerated cosmic rays interacting with surrounding material through processes such as pion decay or inverse Compton scattering. While its precise distance remains uncertain, estimates place it at approximately 10,400 light-years (or $\sim 3.2$ kiloparsecs) if associated with SNR G353.6-0.7 Sagun et al. (2023); Horvath et al. (2023); Char and Biswas (2024); Zhang et al. (2024).
@@
Alternatively, HESS J1731-347 could be a hadron star (HS) or strange star (SS) with an early deconfinement phase transition occurring below twice nuclear saturation density. Such a star would possess a significant quark-gluon plasma core, potentially leading to rapid cooling via the direct Urca (DU) process involving active quarks. However, the two-color superconducting (2SC) phase, characterized by the absence of strange quarks and the formation of a u-d diquark condensate in specific color directions, suppresses rapid cooling and aligns with the observed surface temperature, as derived from a self-consistent calculation of the quark pairing gap within a chirally symmetric relativistic density functional (RDF) model Sagun et al. (2023).
Quark matter can be categorized based on the presence of strange quarks, with the 2SC phase defined by the absence of strange quarks and the CFL phase characterized by Cooper pairs where color and flavor properties are correlated in a one-to-one correspondence among three colors and three flavors Alford (2009). Additional quark pairing configurations, such as gapless 2SC, crystalline color superconductivity, or gapless CFL (gCFL), are also possible Buballa (2005). The observed surface temperature of HESS J1731-347 aligns well with the 2SC phase and a first-order deconfinement phase transition, negating the need for the CFL phase, which would further suppress neutrino emission while still fitting the data Klähn and Fischer (2015). However, a hybrid equation of state (EoS) combining the MDI-APR1 (hadronic) and CFL (quark) EoS, with a phase transition modeled via Maxwell construction, can explain HESS J1731-347’s properties using stable CFL quark matter. In contrast, hybrid models incorporating the CFL MIT Bag model fail to account for the masses of the heaviest observed pulsars Kourmpetis and Moustakidis (2025). Studies further suggest that HESS J1731-347 is either a strange star or a hadron star Rather et al. (2023), with comparisons provided in Figures 1 and 2 of Sagun et al. (2023), Figure 3 of Rather et al. (2023), Figure 4 of Klähn and Fischer (2015), Figure 5 of Brandes and Weise (2024), and Figure 12 of Ayriyan et al. (2025).
@@ 2.4 Stranglet Formation in Astrophysical and Experimental Contexts
In neutron stars, such as HESS J1731-347, the CFL phase’s stability under neutrality constraints supports strangelet formation. In heavy-ion collisions, transient 2SC or CFL phases may produce strangelets, potentially detectable through negative pion radiation or proton excess Rapp et al. (1998); Kourmpetis and Moustakidis (2025). The CFL phase is more favorable for strangelet formation than the 2SC phase due to its inclusion of strange quarks and lower free energy. However, instabilities in mixed phases and external factors like magnetic fields introduce uncertainties. Observational signatures, such as r-mode damping in neutron stars or particle production in collisions, could provide evidence for strangelets Kourmpetis and Moustakidis (2025); Shovkovy (2005).
@@ 3 Future Experimental Goals
3.1 The Cherenkov Telescope Array Observatory for Strangelet Searches
The Cherenkov Telescope Array (CTA) is a next-generation ground-based observatory designed to detect very-high-energy (VHE) gamma rays from 20 GeV to over 300 TeV, establishing it as the most advanced facility for gamma-ray astronomy Abdalla et al. (2021). The CTA aims to explore high-energy astrophysical phenomena, including cosmic ray acceleration, dark matter, and exotic particles like strangelets, particularly in the context of HESS J1731-347 within the TeV gamma-ray range (0.1-100 TeV) Abramowski et al. (2011); Capasso et al. (2017). Unlike optical telescopes, the CTA detects gamma rays indirectly by observing Cherenkov radiation, blue light flashes produced when VHE gamma rays create particle showers in Earth’s atmosphere Cerenkov et al. (1934). Comprising two arrays for full-sky coverage, the Southern Array (CTA-South) at Paranal Observatory, Chile (2,600 m altitude), is optimized for galactic sources. The 1600-hour Galactic Plane Survey (GPS) will map TeV sources, including HESS J1731-347 Sousa et al. (2025).
@@ 3.2 CTA Spectral Line Detection for HESS J1731-347
The CTA offers unparalleled sensitivity to gamma-ray lines in the 0.1-10 TeV range, where its effective area is optimized for detecting the gamma-ray flux of HESS J1731-347. The differential flux at 1 TeV is $\frac{dN}{dE}=(3.0\pm 0.6_{\rm stat}\pm 0.7_{\rm syst})\times 10^{-12}$ cm-2 s-1 TeV-1, with an integral flux above 1 TeV of approximately $3.5\times 10^{-12}$ cm-2 s-1, roughly 3% of the Crab Nebula’s flux Doroshenko et al. (2023); Guo et al. (2018). Strangelets from HESS J1731-347 could produce line-like features through interactions with the interstellar medium (ISM) or exotic decay processes Madsen (1999); Bucciantini et al. (2022). With an energy resolution of approximately 10% at 1 TeV, the CTA can detect lines with fluxes as low as $\sim 10^{-13}$ ph/cm2/s after 100 hours of observation, compared to H.E.S.S.’s limit of $\sim 10^{-12}$ ph/cm2/s Abdalla et al. (2021). For HESS J1731-347, with a flux of $\sim 10^{-12}$ ph/cm2/s at 1 TeV, the CTA can probe contributions below 0.1% of the continuum flux. By fitting the spectrum to a power law ($dN/dE\propto E^{-2.3}$) and searching for Gaussian peaks, the strangelet flux can be constrained as follows:
[…]
Where $n_{H}$ is the average column density, $\sigma_{s}$ is the cross-section, and $F_{\gamma}^{\text{line}}$ is the line photon flux. With an angular resolution of approximately 1 arcminute, the CTA can map HESS J1731-347’s morphology, correlating emission with ISM gas or neutron star remnants. Localized excesses could indicate strangelet production Madsen (1999). The GPS, with $\sim$100 hours of observation, achieves sensitivity to faint signals ($<10^{-13}$ ph/cm2/s), enabling tests of strangelet stability Sousa et al. (2025).
@@ 3.3 CTA Simulations for HESS J1731-347
No studies have explicitly modeled strangelet spectral lines for the CTA in the context of HESS J1731-347 Bauswein et al. (2010); Madsen (1999), as proton-driven $\pi^{0}$ hadronic gamma-ray emission dominates the observed spectrum. However, the CTA’s Key Science Projects (KSPs) for dark matter and axion-like particles provide a framework for simulating strangelet-ISM interactions using the Gamma-ray Astronomy Science Analysis Software (ctools) Knödlseder et al. (2021a) based on GammaLib (A versatile toolbox for scientific analysis of astronomical gamma-ray data) Knödlseder et al. (2021b), and Gammapy (A Python package for gamma-ray astronomy) Donath et al. (2023) to predict spectral signatures Abdalla et al. (2021). To demonstrate the CTA’s potential, we present a chart illustrating the upper limit on a hypothetical strangelet-induced spectral line flux versus gamma-ray energy (0.1-10 TeV), compared to H.E.S.S. limits and HESS J1731-347’s continuum flux.
@@
Figure 1: CTA limitations for strangelet spectral line flux (0.1-10 TeV). The plot depicts the upper CTA limit of strangelet-induced line flux (red), H.E.S.S. limit (magenta), and observed continuum flux (blue) for HESS J1731-347, with a logarithmic y-axis for flux (ph/cm2/s).
@@
Figure 2: Similar to Figure 1, with a logarithmic x-axis, the CTA limit line at $\sim 10^{-13}$ ph/cm2/s, and enhanced Monte Carlo simulation. Simulation results for CTA limits over $\lesssim 10^{-13}$ ph/cm2/s beyond 2 TeV are unsatisfactory, as the HESS J1731-347 cutoff energy began at about 7-10 TeV.
The Figure 1 and 2’s x-axis spans 0.1-10 TeV, where the CTA’s Medium-Sized Telescopes (MSTs) and Small-Sized Telescopes (SSTs) are most sensitive. The y-axis, on a logarithmic scale, represents line flux (ph/cm2/s), with the CTA’s limit at $\sim 10^{-13}$ ph/cm2/s, reflecting its tenfold improvement over H.E.S.S.’s $\sim 10^{-12}$ ph/cm2/s. The continuum flux follows a power law ($dN/dE\propto E^{-2.3}$) Capasso et al. (2017). Theoretical constraints indicate a line flux for HESS J1731-347 of $<10^{-13}$ ph/cm2/s at 1 TeV, with a strangelet flux of $F_{s}<2.5\times 10^{-11}$ ph/cm2/s/sr. The number density is estimated at $<3\times 10^{-22}$ cm-3 for strangelets with a total mass of $<10^{-17}M_{\odot}$ and a production rate of $<10^{-12}$ strangelets/s Bauswein et al. (2010).

### v2

@@ Abstract
Abstract—The hypothesis that supernova remnants are key sources of Galactic cosmic rays gains support from evidence that HESS J1731-347 one of the few Galactic objects capable of accelerating hadronic cosmic rays to TeV energies may harbor an exotic strange quark star rather than a conventional neutron star. This conclusion stems from its unusually low mass and compact radius, which challenge standard neutron star models. If confirmed, such a quark star could generate cosmic rays through the transition from the two-flavor color-superconducting (2SC) phase to the color-flavor-locked (CFL) phase, potentially releasing strangelets, hypothetical strange quark matter (SQM) particles. Detecting these strangelets in cosmic rays would provide groundbreaking evidence for quark matter. The future Cherenkov Telescope Array (CTA), with its unmatched sensitivity and spectral resolution in the very-high-energy (VHE) gamma-ray regime, is uniquely positioned to search for their annihilation or decay signatures. We analyze theoretical predictions for these gamma-ray signals and evaluate CTA’s potential to detect or constrain them. Additionally, we present an in-depth assessment of CTA observations of HESS J1731-347, focusing on spectral features that could confirm strangelet production. A positive detection would not only validate the existence of strange quark stars but also establish a direct link between quark matter and cosmic-ray acceleration, reshaping our understanding of compact objects and high-energy astrophysics.
@@ 1 Introduction
Should strangelets undergo disintegration or annihilation, they could produce distinctive spectral features in the gamma-ray spectrum. The Cherenkov Telescope Array (CTA) Observatory, with its exceptional sensitivity and energy resolution, is ideally suited to search for these signals. As the most advanced facility for gamma-ray astronomy, the CTA is designed to detect very-high-energy (VHE) gamma rays in the energy range of 20 GeV to over 300 TeV. It can probe extreme astrophysical environments, including active galactic nuclei, supernova remnants, and potentially exotic phenomena involving strangelets or dark matter. With superior sensitivity, resolution, and sky coverage compared to existing instruments like MAGIC, VERITAS, and H.E.S.S., the CTA is poised to investigate high-energy astrophysical phenomena, such as possible signatures of strangelets in the context of HESS J1731-347. This document summarizes the theoretical predictions for strangelet-related gamma-ray signatures, their detectability with the CTA, and their implications for constraining the properties of strangelets and HESS J1731-347.
@@ 2.1 Strange Quark Matter and Current Searches
The concept of SQM was first proposed in the 1980s as a distinct state of hadronic matter, coexisting alongside conventional nuclear matter Bodmer (1971); Witten (1984); Farhi and Jaffe (1984). The mass of strangelets within compact stars can range from a minimum stable mass of approximately $10sA$ Chodos et al. (1974) to as high as $A\simeq 10^{57}$, equivalent to a solar mass of approximately $2\times 10^{30}$ kg Haensel et al. (1986); Madsen (2005). SQM could have formed in the early universe shortly after the Big Bang Chodos et al. (1974), potentially contributing to baryonic dark matter Atreya et al. (2014). It may reside within the cores of neutron stars or exist as ‘‘strange quark stars,’’ which could be either purely composed of quarks Haensel et al. (1986); Alcock et al. (1986) or a mixture of hadrons and quarks Drago et al. (2016); Drago and Pagliara (2016); Drago et al. (2014). Strangelets could also be accelerated in the ergosphere of black holes Bianchi et al. (2020) or produced in collisions between binary neutron stars Bucciantini et al. (2022); Drago and Pagliara (2018); Wiktorowicz et al. (2017), potentially reaching Earth De Rújula and Glashow (1984).
[…]
@@ 2.2 HESS J1731-347 and Strong First-Order (Deconfinement) Phase Transition
The High Energy Stereoscopic System (H.E.S.S.), an array of imaging atmospheric Cherenkov telescopes in Namibia, discovered the VHE gamma-ray source HESS J1731-347. Located in the constellation Scorpius near the Galactic plane, it has celestial coordinates of right ascension (RA) 17 h 31 m, declination (Dec) $34^{\circ}47^{\prime}$, and galactic coordinates of longitude $l\simeq 353.6^{\circ}$ (near the galactic center) and latitude $b\simeq-0.7^{\circ}$ (slightly below the galactic plane). Its spatial coincidence with the shell-type supernova remnant (SNR) G353.6-0.7 suggests a potential association. The gamma-ray emission, observed at energies above 100 GeV, may result from accelerated cosmic rays interacting with surrounding material through processes such as pion decay or inverse Compton scattering. While its precise distance remains uncertain, estimates place it at approximately 10400 light-years (or $\sim$3.2 kiloparsecs) if associated with SNR G353.6-0.7 Sagun et al. (2023); Horvath et al. (2023); Char and Biswas ; Zhang et al. (2024).
@@
Alternatively, HESS J1731-347 could be a hadron star (HS) or strange star (SS) with an early deconfinement phase transition occurring below twice nuclear saturation density. Such a star would possess a significant quark–gluon plasma core, potentially leading to rapid cooling via the direct Urca (DU) process involving active quarks. However, the two-color superconducting (2SC) phase, characterized by the absence of strange quarks and the formation of a u-d diquark condensate in specific color directions, suppresses rapid cooling and aligns with the observed surface temperature, as derived from a self-consistent calculation of the quark pairing gap within a chirally symmetric relativistic density functional (RDF) model Sagun et al. (2023).
Quark matter can be categorized based on the presence of strange quarks, with the 2SC phase defined by the absence of strange quarks and the CFL phase characterized by Cooper pairs where color and flavor properties are correlated in a one-to-one correspondence among three colors and three flavors Alford (2009). Additional quark pairing configurations, such as gapless 2SC, crystalline color superconductivity, or gapless CFL (gCFL), are also possible Buballa (2005). The observed surface temperature of HESS J1731-347 aligns well with the 2SC phase and a first-order deconfinement phase transition, negating the need for the CFL phase, which would further suppress neutrino emission while still fitting the data Klähn and Fischer (2015). However, a hybrid equation of state (EoS) combining the MDI-APR1 (hadronic) and CFL (quark) EoS, with a phase transition modeled via Maxwell construction, can explain HESS J1731-347’s properties using stable CFL quark matter. In contrast, hybrid models incorporating the CFL MIT Bag model fail to account for the masses of the heaviest observed pulsars Kourmpetis et al. (2025). Studies further suggest that HESS J1731-347 is either a strange star or a hadron star Rather et al. (2023), with comparisons provided in Figs 1 and 2 of Sagun et al. (2023), Fig 3 of Rather et al. (2023), Fig 4 of Klähn and Fischer (2015), Fig 5 of Brandes and Weise (2024), and Fig 12 of Ayriyan et al. (2025).
@@ 2.4 Stranglet Formation in Astrophysical and Experimental Contexts
In neutron stars, such as HESS J1731-347, the CFL phase’s stability under neutrality constraints supports strangelet formation. In heavy-ion collisions, transient 2SC or CFL phases may produce strangelets, potentially detectable through negative pion radiation or proton excess Rapp et al. (1998); Kourmpetis et al. (2025). The CFL phase is more favorable for strangelet formation than the 2SC phase due to its inclusion of strange quarks and lower free energy. However, instabilities in mixed phases and external factors like magnetic fields introduce uncertainties. Observational signatures, such as r-mode damping in neutron stars or particle production in collisions, could provide evidence for strangelets Kourmpetis et al. (2025); Shovkovy (2005).
@@ 3 Future Experimental Goals
3.1 The CTA Observatory for Strangelet Searches
The CTA is a next-generation ground-based observatory designed to detect VHE gamma rays from 20 GeV to over 300 TeV, establishing it as the most advanced facility for gamma-ray astronomy Abdalla et al. (2021). The CTA aims to explore high-energy astrophysical phenomena, including cosmic ray acceleration, dark matter, and exotic particles like strangelets, particularly in the context of HESS J1731-347 within the TeV gamma-ray range (0.1–100 TeV) Abramowski et al. (2011); Capasso et al. (2017). Unlike optical telescopes, the CTA detects gamma rays indirectly by observing Cherenkov radiation, blue light flashes produced when VHE gamma rays create particle showers in Earth’s atmosphere Cerenkov et al. (1934). Comprising two arrays for full-sky coverage, the Southern Array (CTA-South) at Paranal Observatory, Chile (2600 m altitude), is optimized for galactic sources. The 1600-hour Galactic Plane Survey (GPS) will map TeV sources, including HESS J1731-347 Sousa et al. (2025).
The process $SS\to\gamma\gamma$ describes a theoretical scenario in which two strangelets ($S$), hypothetical particles made of SQM, annihilate to produce two gamma-ray photons. The energy of each gamma-ray photon is given by $E_{\gamma}=m_{S}c^{2}$, where $m_{S}$ is the rest mass of the strangelet and $c$ is the speed of light. This detection is facilitated by observing Cherenkov radiation produced by air showers in Earth’s atmosphere Acharya et al. (2013); Bernlöhr et al. (2013).
@@ 3.2 CTA Spectral Line Detection for HESS J1731-347
The CTA offers unparalleled sensitivity to gamma-ray lines in the 0.1–10 TeV range, where its effective area is optimized for detecting the gamma-ray flux of HESS J1731-347. The differential flux at 1 TeV is $\frac{dN}{dE}=(3.0\pm 0.6_{\rm stat}\pm 0.7_{\rm syst})\times 10^{-12}$ cm-2 s-1 TeV-1, with an integral flux above 1 TeV of approximately $3.5\times 10^{-12}$ cm-2 s-1, roughly 3% of the Crab Nebula’s flux Doroshenko et al. (2023); Guo et al. (2018). Strangelets from HESS J1731-347 could produce line-like features through interactions with the interstellar medium (ISM) or exotic decay processes Madsen (1999); Bucciantini et al. (2022). With an energy resolution of approximately 10% at 1 TeV, the CTA can detect lines with fluxes as low as $\sim$10-13 ph/cm2/s after 100 hours of observation, compared to H.E.S.S.’s limit of $\sim$10-12 ph/cm2/s Abdalla et al. (2021). For HESS J1731-347, with a flux of $\sim$10-12 ph/cm2/s at 1 TeV, the CTA can probe contributions below 0.1% of the continuum flux. By fitting the spectrum to a power law ($dN/dE\propto E^{-2.3}$) and searching for Gaussian peaks, the strangelet flux can be constrained as follows:
$F_{s}<\frac{F_{\gamma}^{\text{line}}}{n_{H}\cdot\sigma_{s}}\approx\frac{10^{-13}}{100\cdot 4\times 10^{-26}}\approx 2.5\times 10^{-11}\,\text{ph/cm}^{2}/\text{s/sr}.$ (1)
Where $n_{H}$ is the average column density, $\sigma_{s}$ is the cross-section, and $F_{\gamma}^{\text{line}}$ is the line photon flux. With an angular resolution of approximately 1 arcminute, the CTA can map HESS J1731-347’s morphology, correlating emission with ISM gas or neutron star remnants. Localized excesses could indicate strangelet production Madsen (1999). The GPS, with $\sim$100 hours of observation, achieves sensitivity to faint signals ($10^{-13}$ ph/cm2/s), enabling tests of strangelet stability Sousa et al. (2025).
@@ 3.3 CTA Simulations for HESS J1731-347
No studies have explicitly modeled strangelet spectral lines for the CTA in the context of HESS J1731-347 Bauswein et al. (2010); Madsen (1999), as proton-driven $\pi^{0}$ hadronic gamma-ray emission dominates the observed spectrum. However, the CTA’s Key Science Projects (KSPs) for dark matter and axion-like particles provide a framework for simulating strangelet-ISM interactions using the Gamma-ray Astronomy Science Analysis Software (ctools) Knödlseder et al. (2021a) based on GammaLib (A versatile toolbox for scientific analysis of astronomical gamma-ray data) Knödlseder et al. (2021b), and Gammapy (A Python package for gamma-ray astronomy) Donath et al. (2023) to predict spectral signatures Abdalla et al. (2021). To demonstrate the CTA’s potential, we present a chart illustrating the upper limit on a hypothetical strangelet-induced spectral line flux versus gamma-ray energy (0.1–10 TeV), compared to H.E.S.S. limits and HESS J1731-347’s continuum flux.
@@
Figure 1: CTA limitations for strangelet spectral line flux (0.1–10 TeV). The plot depicts the upper CTA limit of strangelet-induced line flux (red), H.E.S.S. limit (magenta), and observed continuum flux (blue with error-band) for HESS J1731-347, with a logarithmic y-axis for flux (ph/cm2/s).
@@
Figure 2: Similar to Fig 1, with a logarithmic x-axis, the CTA limit line at $\sim$10-13 ph/cm2/s, and enhanced Monte Carlo simulation. Simulation results for CTA limits over $\lesssim$10-13 ph/cm2/s beyond 2 TeV are unsatisfactory, as the HESS J1731-347 cutoff energy began at about 7–10 TeV.
The Figs 1 and 2’s x-axis spans 0.1–10 TeV, where the CTA’s Medium-Sized Telescopes (MSTs) and Small-Sized Telescopes (SSTs) are most sensitive. The y-axis, on a logarithmic scale, represents line flux (ph/cm2/s), with the CTA’s limit at $\sim$10-13 ph/cm2/s, reflecting its tenfold improvement over H.E.S.S.’s $\sim$10-12 ph/cm2/s. The continuum flux follows a power law ($dN/dE\propto E^{-2.3}$) Capasso et al. (2017). Theoretical constraints indicate a line flux for HESS J1731-347 of $10^{-13}$ ph/cm2/s at 1 TeV, with a strangelet flux of $F_{s}<2.5\times 10^{-11}$ ph/cm2/s/sr. The number density is estimated at $3\times 10^{-22}$ cm-3 for strangelets with a total mass of $10^{-17}M_{\odot}$ and a production rate of $10^{-12}$ strangelets/s Bauswein et al. (2010).